# CLIP-QDA: An Explainable Concept Bottleneck Model

### v1

@@ 5 Conclusion
In this paper, we introduce a modeling approach for the embedding space of CLIP, a foundation model trained on a vast dataset. Our observations reveal that CLIP can organize information in a distribution that exhibits similarities with a mixture of Gaussians. Building upon this insight, we develop an adapted concept bottleneck model that demonstrates competitive performance along with transparency. While the model that we have presented offers the advantage of simplicity and a limited number of parameters, it does encounter challenges when dealing with a broader and more ambiguous set of concepts. As a suggestion for future research, we propose to extend this modeling approach to incorporate other priors, such as Laplacian distributions, and to explore more complex models, including those with multiple components, i.e., using more than one Gaussian to describe a class for example. Additionally, our research is centered around a specific embedding space (CLIP scores). Exploring similar work on other latent spaces, particularly those associated with multimodal foundation models, could be valuable to determine if similar patterns exist in those spaces.
@@ Appendix A Appendix
A.1 Set of concepts
Inspired by Yang et al. (2023), we use large language models to provide concept sets. Concretely, we use GPT-3 (Brown et al., 2020) with the following preprompt: “In this task, you have to give visual descriptions that describe an image. Respond as a list. Each item being a word.” Then, we generate the sets of words by the following prompt: “What are [N] useful visual descriptors to distinguish a [class] in a photo?” By doing so, we generated 5 concepts per class, presented in Table 2.
Table 2: List of concepts used.
@@ A.1 Set of concepts
’aeroplane’:[’Winged, Jet engines, Tail fin, Fuselage, Landing gear’],’bicycle’:[’Two-wheeled, Pedals, Handlebars, Frame, Chain-driven’],’bird’:[’Feathery, Beaked, Wingspread, Perched, Avian’],’bottle’:[’Glass or plastic, Cylindrical, Necked, Cap or cork, Transparent or translucent’],’bus’:[’Large, Rectangular, Windows, Wheels, Multi-doored’],’cat’:[’Furry, Whiskered, Pointy-eared, Slitted-eyed, Four-legged’],’car’:[’Metallic, Four-wheeled, Headlights, Windshield, License plate’],’dog’:[’Snout, Wagging-tailed, Snout-nosed, Floppy-eared, Tail-wagging’],’cow’:[’Bovine, Hooved, Horned (in some cases), Spotted or solid-colored, Grazing (if in a field)’],’horse’:[’Equine, Hooved, Maned, Tailed, Galloping (if in motion)’],’motorbike’:[’Two-wheeled , Engine, Handlebars , Exhaust, Saddle or seat’],’person’:[’Human, Facial features, Limbs (arms and legs), Clothing, Posture or stance’],’pottedplant’:[’Potted, Green, Leaves, Soil or potting mix, Container or pot’],’sheep’:[’Woolly, Hooved, Grazing, Herded (if in a group), White or colored fleece’],’train’:[’Locomotive, Railcars, Tracks, Wheels , Carriages’],’tvmonitor’:[’Screen, Rectangular , Frame or bezel, Stand or wall-mounted, Displaying an image’]
Monumai
@@
’Baroque’:[’Ornate, Elaborate sculptures, Intricate details, Curved or asymmetrical design, Historical or aged appearance’],’Gothic’:[’Pointed arches, Ribbed vaults, Flying buttresses, Stained glass windows, Tall spires or towers’],’Hispanic muslim’:[’Mudejar style, Intricate geometric patterns, Horseshoe arches, Decorative tilework (azulejos), Islamic-inspired motifs’],’Rennaissance’:[’Classical proportions, Symmetrical design, Columns and pilasters, Human statues and sculptures, Dome or dome-like structures’]
MIT scenes
@@
’Store’:[’Building or structure, Signage or banners, Glass windows or doors, Displayed products or merchandise, People entering or exiting (if applicable)’],’Home’:[’Residential, Roofed, Windows, Landscaping or yard, Front entrance or door’],’Public space’:[’Open area, Crowds (if people are present), Benches or seating, Pathways or walkways, Architectural features (e.g., buildings, statues)’],’Leisure’:[’Recreational, Play equipment (if applicable), Greenery or landscaping, Picnic tables or seating, Relaxing atmosphere’],’Working space’:[’Office equipment (e.g., desks, computers), Task-oriented, Office chairs, Organized or structured, People working (if applicable)’]
Cats/Dogs/Cars
@@
’Cat’:[’Furry, Whiskered, Pointy-eared, Slitted-eyed, Four-legged’],’Car’:[’Metallic, Four-wheeled, Headlights, Windshield, License plate’],’Dog’:[’Snout, Wagging-tailed, Snout-nosed, Floppy-eared, Tail-wagging’]
A.2 Closed-form solution of counterfactuals for CLIP QDA
@@
$\min\|{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}\|^{2}~{}~{}\textnormal{s.t.}\quad\frac{p_{c_{\overline{h}}}}{(2\pi)^{N/2}|{\bm{\Sigma}}_{c_{\overline{h}}}|^{\frac{1}{2}}}e^{-\frac{1}{2}({\bm{z}}+{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}-{\bm{\mu}}_{c_{\overline{h}}})^{T}{\bm{\Sigma}}_{c_{\overline{h}}}^{-1}({\bm{z}}+{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}-{\bm{\mu}}_{c_{\overline{h}}})}\leqslant\frac{p_{c_{h}}}{(2\pi)^{N/2}|{\bm{\Sigma}}_{c_{h}}|^{\frac{1}{2}}}e^{-\frac{1}{2}({\bm{z}}+{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}-{\bm{\mu}}_{{c_{h}}})^{T}{\bm{\Sigma}}_{c_{h}}^{-1}({\bm{z}}+{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}-{\bm{\mu}}_{{c_{h}}})}.$ (9)
@@
The third equation of 12 indicates whether the inequality condition is saturated or not. If it is not saturated ($\lambda=0$), then the label $c_{\overline{h}}$ is already achieved for ${\bm{z}}$, resulting in a solution of $\epsilon^{j}_{s,*}=0$. This being impossible by construction, we only focus on the case where $\lambda\neq 0$.
@@
Considering 13, the validity of the results is conditioned by the sign $s$ and the condition of the magnitude of $\epsilon^{j}_{s,*}$. Then, the final result of the problem 5 is either $\emptyset$, $b_{1}$ or $b_{2}$ depending on the conditions:
@@
To expand problem 5 to multiclass classification $C>2$, we consider it as a succession of $C-1$ binary classifications between each $i^{\prime}\neq c_{h}$ and $c_{h}$. Given a concept index $j$ and a sign $s$, if we denote the solutions of theses problems as the set $\{{\bm{\epsilon}}^{j}_{+,*,1},...,{\bm{\epsilon}}^{j}_{s,*,c_{h}-1},{\bm{\epsilon}}^{j}_{s,*,c_{h}+1},...,{\bm{\epsilon}}^{j}_{s,*,C}\}$, the final solution ${\bm{\epsilon}}^{j}_{s,*}$ is, if it exists, the value of minimal magnitude among this set.
Examples of explanations based on this metric are given in Sections 4.4 and A.4.
A.3 Local explanation time
In Table 3, we display the amount of time taken to produce a local explanation for each image of the test set of PASCAL-Part and Cats/Dogs/Cars.
@@ A.3 Local explanation time
Method ours (local) LIME SHAP
@@
A.4 Additional samples
We display here additional samples of both local and global explanations from the PASCAL-part dataset. For clarity purposes, we used these examples with the following set of 10 concepts: [“Winged”, “Pedals”, “Metallic”, “Clothing”, “Wheels”, “Hooved”, “Feathery”, “Green”, “Tracks”, “Furry”].
(a) Input image
(b) Probabilities classifier Only the top 5 classes are displayed
(c) Local explanation Only the top 5 classes are displayed
(d) Global explanation associated with the two highest probabilities. Here “aeroplane” and “motorbike”
Figure 12: Example of explanation on the PASCAL-Part dataset
(a) Input image
(b) Probabilities classifier Only the top 5 classes are displayed
(c) Local explanation Only the top 5 classes are displayed
(d) Global explanation associated with the two highest probabilities. Here “potted plant” and “bottle”
Figure 13: Example of explanation on the PASCAL-Part dataset
(a) Input image
(b) Probabilities classifier Only the top 5 classes are displayed
(c) Local explanation Only the top 5 classes are displayed
(d) Global explanation associated with the two highest probabilities. Here “person” and “car”
Figure 14: Example of explanation on the PASCAL-Part dataset
(a) Input image
(b) Probabilities classifier Only the top 5 classes are displayed
(c) Local explanation Only the top 5 classes are displayed
(d) Global explanation associated with the two highest probabilities. Here “person” and “horse”
Figure 15: Example of explanation on the PASCAL-Part dataset

### v2

@@ 5 Conclusion
In this paper, we introduce a modeling approach for the embedding space of CLIP, a foundation model trained on a vast dataset. Our observations reveal that CLIP can organize information in a distribution that exhibits similarities with a mixture of Gaussians. Building upon this insight, we develop an adapted concept bottleneck model that demonstrates competitive performance along with transparency. While the model that we have presented offers the advantage of simplicity and a limited number of parameters, it does encounter challenges when dealing with a broader and more ambiguous set of concepts. As a suggestion for future research, we propose to extend this modeling approach to incorporate other priors, such as Laplacian distributions, and to explore more complex models, including those with multiple components, i.e., using more than one Gaussian to describe a class for example. Another avenue to potentially enhance the performance of our model is to explore guiding the latent space towards a Gaussian distribution. Additionally, our research is centered around a specific embedding space (CLIP scores). Exploring similar work on other latent spaces, particularly those associated with multimodal foundation models, could be valuable to determine if similar patterns exist in those spaces.
@@ Appendix A Appendix
A.1 Implementation details
GradCAM.
To compute GradCAM explanations, we applied the method to the $21^{th}$ block of the image encoder in CLIP-QDA.
LIME.
The LIME image-level explanations are generated using the superpixel approach defined in (Ribeiro et al., 2016). This method involves treating the input as a set of superpixels obtained through a segmentation process. LIME is then applied, considering these superpixels as binary values corresponding to whether they contribute to displaying the related zone. Following the original implementation, we use quickshift clustering to segment the image. We display the top 10 superpixels that exert the most influence on the prediction, highlighting superpixels in blue that have a negative impact on the label and superpixels in red that have a positive impact on the label.
SHAP.
Following the advised method from the SHAP library, the image-level explanation from SHAP is computed using the partition explainer, which creates from the input a subset of images by applying a set of perturbations. In all our experiments, the perturbations are a blur of kernel (128,128).
CLIP-LIME.
For the CLIP-LIME, we consider inputs as a set of tabular values. Then, each CLIP-LIME explanation is obtained by running the lime tabular explainer provided by the author’s repo with $num~{}features=100000$, $num~{}samples=1000$, and $batchsize=10$.
CLIP-SHAP.
For the CLIP-SHAP, we consider inputs as a set of tabular values. Then, each CLIP-SHAP explanation is obtained by running the shap kernel explainer provided by the author’s repo with a background of $5$ samples using k-means.
LaBo.
For LaBo, we train a LaBo classifier by using Adam optimizer with a learning rate of $15e-$ 1 a weight decay of 00, and a batch size of $8192$. The resulting explanations follow the resulting weight matrix.
Yan et al.
For the Yan et al. method, we train a linear classifier by using Adam optimizer with a learning rate of $35e-$ 3 a weight decay of $41e-$ 4, and a batch size of $512$. The sample-wise (concept level) explanation results from the product between the concept score and its associated weight. The dataset-wise explanation results from the weights alone.
A.2 Cats/Dogs/Cars
The Cats/Dogs/Cars dataset is available in https://huggingface.co/datasets/RemiKaz/Cats_Dogs_Cars. To create this dataset, we combined two well-known datasets: the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). We then filtered the dataset to include only images featuring black and white animals and cars, leading to six different categories. The primary objective behind assembling this dataset is to facilitate research in scenarios characterized by significant data bias, such as the classification of white cats when the training data predominantly consists of images of white dogs and black cats.
A.3 Set of concepts
Inspired by Yang et al. (2023), we use large language models to provide concept sets. Concretely, we use GPT-3 (Brown et al., 2020) with the following preprompt: “In this task, you have to give visual descriptions that describe an image. Respond as a list. Each item being a word.” Then, we generate the sets of words by the following prompt: “What are [N] useful visual descriptors to distinguish a [class] in a photo?” By doing so, we generated 5 concepts per class, presented in Table 6. An additional ordering of concepts by subcategories is available in Table 7.
Table 6: List of concepts used, ordered by classes of interest.
@@ A.3 Set of concepts
PASCAL-Part ’aeroplane’:[Winged, Jet engines, Tail fin, Fuselage, Landing gear],’bicycle’:[Two-wheeled, Pedals, Handlebars, Frame, Chain-driven],’bird’:[Feathery, Beaked, Wingspread, Perched, Avian],’bottle’:[Glass or plastic, Cylindrical, Necked, Cap or cork, Transparent or translucent],’bus’:[Large, Rectangular, Windows, Wheels, Multi-doored],’cat’:[Furry, Whiskered, Pointy-eared, Slitted-eyed, Four-legged],’car’:[Metallic, Four-wheeled, Headlights, Windshield, License plate],’dog’:[Snout, Wagging-tailed, Snout-nosed, Floppy-eared, Tail-wagging],’cow’:[Bovine, Hooved, Horned (in some cases), Spotted or solid-colored, Grazing (if in a field)],’horse’:[Equine, Hooved, Maned, Tailed, Galloping (if in motion)],’motorbike’:[Two-wheeled , Engine, Handlebars , Exhaust, Saddle or seat],’person’:[Human, Facial features, Limbs (arms and legs), Clothing, Posture or stance],’potted plant’:[Potted, Green, Leaves, Soil or potting mix, Container or pot],’sheep’:[Woolly, Hooved, Grazing, Herded (if in a group), White or colored fleece],’train’:[Locomotive, Railcars, Tracks, Wheels , Carriages],’tvmonitor’:[Screen, Rectangular , Frame or bezel, Stand or wall-mounted, Displaying an image]
@@
MonuMAI ’Baroque’:[Ornate, Elaborate sculptures, Intricate details, Curved or asymmetrical design, Historical or aged appearance],’Gothic’:[Pointed arches, Ribbed vaults, Flying buttresses, Stained glass windows, Tall spires or towers],’Hispanic muslim’:[Mudejar style, Intricate geometric patterns, Horseshoe arches, Decorative tilework (azulejos), Islamic-inspired motifs],’Rennaissance’:[Classical proportions, Symmetrical design, Columns and pilasters, Human statues and sculptures, Dome or dome-like structures]
@@
MIT scenes ’Store’:[Building or structure, Signage or banners, Glass windows or doors, Displayed products or merchandise, People entering or exiting (if applicable)],’Home’:[Residential, Roofed, Windows, Landscaping or yard, Front entrance or door],’Public space’:[Open area, Crowds (if people are present), Benches or seating, Pathways or walkways, Architectural features (e.g., buildings, statues)],’Leisure’:[Recreational, Play equipment (if applicable), Greenery or landscaping, Picnic tables or seating, Relaxing atmosphere],’Working space’:[Office equipment (e.g., desks, computers), Task-oriented, Office chairs, Organized or structured, People working (if applicable)]
@@
Cats/Dogs/Cars ’Cat’:[Furry, Whiskered, Pointy-eared, Slitted-eyed, Four-legged],’Car’:[Metallic, Four-wheeled, Headlights, Windshield, License plate],’Dog’:[Snout, Wagging-tailed, Snout-nosed, Floppy-eared, Tail-wagging]
Table 7: List of concepts used, ordered by subcategories.
Dataset Set of concepts
PASCAL-Part ’Objects’:[Screen, Rectangular, Frame or bezel, Stand or wall-mounted, Displaying an image, Glass or plastic, Cylindrical, Necked, Cap or cork, Transparent or translucent, Potted, Green, Leaves, Soil or potting mix, Container or pot], ’Transportation-related’:[Two-wheeled, Engine, Handlebars, Exhaust, Saddle or seat, Four-wheeled, Headlights, Windshield, License plate, Wheels, Multi-doored, Locomotive, Railcars, Tracks, Carriages, Chain-driven],’Aircraft-related’:[Winged, Jet engines, Tail fin, Fuselage, Landing gear],’Building/Structure-related’:[Large, Rectangular, Windows] ’Human Characteristics’:[Human, Facial features, Limbs (arms and legs), Clothing, Posture or stance],’Avian Characteristics’:[Feathery, Beaked, Wingspread, Perched, Avian],’Animal Characteristics’:[Furry, Whiskered, Pointy-eared, Slitted-eyed, Four-legged, Equine, Hooved, Maned, Tailed, Galloping (if in motion), Snout, Wagging-tailed, Snout-nosed, Floppy-eared, Tail-wagging, Woolly, Grazing, Herded (if in a group), White or colored fleece, Bovine, Horned (in some cases), Spotted or solid-colored]
MonuMAI ’Architectural Styles and Elements’:[Mudejar style, Intricate geometric patterns, Horseshoe arches, Decorative tilework (azulejos), Islamic-inspired motifs, Stained glass windows, Pointed arches, Ribbed vaults],’Artistic Details and Features’:[Ornate, Elaborate sculptures, Intricate details, Curved or asymmetrical design, Human statues and sculptures, Historical or aged appearance],’Architectural Components’:[Flying buttresses, Classical proportions, Symmetrical design, Columns and pilasters, Dome or dome-like structures, Tall spires or towers]
MIT scenes ’Work Environment’:[Office equipment (e.g., desks, computers), Task-oriented, Office chairs, Organized or structured, People working (if applicable)], ’Leisure Environment’:[Play equipment (if applicable), Greenery or landscaping, Picnic tables or seating, Relaxing atmosphere, Recreational],’Community Environment’:[Roofed, Windows, Landscaping or yard, Front entrance or door, Building or structure, Signage or banners, Glass windows or doors, Displayed products or merchandise, People entering or exiting (if applicable), Open area, Crowds (if people are present), Benches or seating, Pathways or walkways, Architectural features (e.g., buildings, statues), Residential]
Cats/Dogs/Cars ’Vehicle Features’:[License plate, Headlights, Windshield, Four-wheeled], ’Animal Features’:[Wagging-tailed, Tail-wagging, Snout, Whiskered, Pointy-eared, Slitted-eyed, Floppy-eared, Furry, Snout-nosed, Four-legged],’Material/Texture’:[Metallic]
A.4 Study of images according to their position in the latent space
In addition to the CLIP-QDAglobal method, we introduce another approach leveraging multivariate Gaussian modeling to generate dataset-wise visual explanations for our model. Specifically, for a given class $c$, we compute the Mahalanobis distance with respect to $\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{c},{\bm{\Sigma}}_{c})$ for each sample’s latent space projection ${\bm{z}}$ labeled as $c$ within a dataset of interest:
$d_{M}({\bm{z}},{\bm{\mu}}_{c},{\bm{\Sigma}}_{c})=\sqrt{({\bm{z}}-{\bm{\mu}}_{c})^{\top}{\bm{\Sigma}}_{c}^{-1}({\bm{z}}-{\bm{\mu}}_{c})}.$ (9)
Then, we plot in Figures 15, 16, 17 and 18, the closest and farthest images with the dataset, according to this distance for PascalPART and MonuMAI.
(a) “Aeroplane”
(b) ‘Bicycle”
(c) “Bird”
(d) “Bottle”
(e) “Bus”
(f) “Car”
(g) “Cat”
(h) “Cow”
(i) “Dog”
(j) “Horse”
(k) “Motorbike”
(l) “Person”
(m) “Potted plant”
(n) “Sheep”
(o) “Train”
(p) “TV monitor”
Figure 15: Images of minimal distance for each classes of PascalPART
(a) “Aeroplane”
(b) ‘Bicycle”
(c) “Bird”
(d) “Bottle”
(e) “Bus”
(f) “Car”
(g) “Cat”
(h) “Cow”
(i) “Dog”
(j) “Horse”
(k) “Motorbike”
(l) “Person”
(m) “Potted plant”
(n) “Sheep”
(o) “Train”
(p) “TV monitor”
Figure 16: Images of maximal distance for each classes of PascalPART
(a) “Baroque”
(b) ‘Gothic”
(c) “Hispanic-Muslim”
(d) “Renaissance”
Figure 17: Images of minimal distance for each classes of MonuMAI
(a) “Baroque”
(b) ‘Gothic”
(c) “Hispanic-Muslim”
(d) “Renaissance”
Figure 18: Images of maximal distance for each classes of MonuMAI
For PascalPART, the closest images tend to be predominantly close-up shots of the object in question. However, this occasionally presents challenges due to the inherent pixelation of the images, leading to somewhat counterintuitive results. Conversely, the furthest images are typically those where the object is distant. Another scenario of failure arises when there are two classes present within the image, which complicates the retrieval process. On the other hand, for MonumAI, the closest images appear to be the simplest ones, without sign of disturbance. In contrast, the farthest images may include extraneous objects, such as signs.
A.5 Closed-form solution of counterfactuals for CLIP QDA
@@
$\min\|{\bm{\epsilon}}^{j}_{{s}}\|^{2}~{}~{}\textnormal{s.t.}\quad\frac{p_{c_{\overline{h}}}}{(2\pi)^{N/2}|{\bm{\Sigma}}_{c_{\overline{h}}}|^{\frac{1}{2}}}e^{-\frac{1}{2}({\bm{z}}+{\bm{\epsilon}}^{j}_{{s}}-{\bm{\mu}}_{c_{\overline{h}}})^{T}{\bm{\Sigma}}_{c_{\overline{h}}}^{-1}({\bm{z}}+{\bm{\epsilon}}^{j}_{{s}}-{\bm{\mu}}_{c_{\overline{h}}})}\leqslant\frac{p_{c_{h}}}{(2\pi)^{N/2}|{\bm{\Sigma}}_{c_{h}}|^{\frac{1}{2}}}e^{-\frac{1}{2}({\bm{z}}+{\bm{\epsilon}}^{j}_{{s}}-{\bm{\mu}}_{{c_{h}}})^{T}{\bm{\Sigma}}_{c_{h}}^{-1}({\bm{z}}+{\bm{\epsilon}}^{j}_{{s}}-{\bm{\mu}}_{{c_{h}}})}.$ (12)
@@
The third equation of 15 indicates whether the inequality condition is saturated or not. If it is not saturated ($\lambda=0$), then the label $c_{\overline{h}}$ is already achieved for ${\bm{z}}$, resulting in a solution of $\epsilon^{j}_{s,*}=0$. This being impossible by construction, we only focus on the case where $\lambda\neq 0$.
@@
Considering 16, the validity of the results is conditioned by the sign $s$ and the condition of the magnitude of $\epsilon^{j}_{s,*}$. Then, the final result of the problem 5 is either $\emptyset$, $b_{1}$ or $b_{2}$ depending on the conditions:
@@
To expand problem 5 to multiclass classification $C>2$, we consider it as a succession of $C-1$ binary classifications between each $i^{\prime}\neq c_{h}$ and $c_{h}$. Given a concept index $j$ and a sign $s$, if we denote the solutions of theses problems as the set $\{{\bm{\epsilon}}^{j}_{+,*,1},...,{\bm{\epsilon}}^{j}_{s,*,c_{h}-1},{\bm{\epsilon}}^{j}_{s,*,c_{h}+1},...,{\bm{\epsilon}}^{j}_{s,*,C}\}$, the final solution ${\bm{\epsilon}}^{j}_{s,*}$ is if it exists, the value of minimal magnitude among this set.
Examples of explanations based on this metric are given in Sections 4.5 and A.7.
A.6 Sample-wise explanation time
In Table 8, we display the amount of time taken to produce a sample-wise explanation for each image of the test set of PASCAL-Part and Cats/Dogs/Cars.
@@ A.6 Sample-wise explanation time
Method CLIP-QDAlocal CLIP-LIME CLIP-SHAP
@@
A.7 Additional samples
We display here additional samples of both sample-wise and dataset-wise explanations from the PASCAL-Part and Cats/Dogs/Cars dataset.
(a) Input image (b) GradCAM Explanation (c) LIME Explanation (d) SHAP explanation
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 19: Sample-wise explanations (image level). The first two examples come from the PascalPART dataset, and the last two samples come from the Cats/Dogs/Cars dataset in the biased setup. Note that the classifier mislabeled the $2^{nd}$ example as “car” and the $4^{th}$ example as “car”.
{adjustwidth}
-2cm2cm (a) Input image (b) CLIP-QDAlocal Explanation (c) Yan et al. (sample) explanation (d) CLIP-LIME explanation Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 20: Sample-wise explanations (concept level). The first two examples come from the PascalPART dataset, and the last two samples come from the Cats/Dogs/Cars dataset in the biased setup. Note that the classifier mislabeled the $2^{nd}$ example as “car” and the $4^{th}$ example as “car”.
{adjustwidth}
-2cm2cm (a) Input image (b) CLIP-SHAP Explanation Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 21: Sample-wise explanations (concept level). The first two examples come from the PascalPART dataset, and the last two samples come from the Cats/Dogs/Cars dataset in the biased setup. Note that the classifier mislabeled the $2^{nd}$ example as “car” and the $4^{th}$ example as “car”.
{adjustwidth}
-2cm2cm (a) Input image (b) CLIP-QDAglobal Explanation (c) LaBo explanation (d) Yan et al. (dataset) explanation Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 22: Dataset-wise explanations. The first two examples come from the PascalPART dataset, and the last two samples come from the Cats/Dogs/Cars dataset in the biased setup. Note that the classifier mislabeled the $2^{nd}$ example as “car” and the $4^{th}$ example as “car”.